OpenStudy (anonymous):

There are 2 liquids, one 10%, one 20%. How much of each mixture would you mix to get 800mL of a 15% liquid?

OpenStudy (mertsj):

x= amount of 10% 800-x= amount of 20% .10x+.20(800-x)=.15(800)

OpenStudy (anonymous):

so x=-11840

OpenStudy (mertsj):

That cannot be.

OpenStudy (mertsj):

\[.10x+160-.20x=120\]

OpenStudy (mertsj):

\[-.10x=-40\]

OpenStudy (mertsj):

\[x=400\]

OpenStudy (anonymous):

i got 400 mL of 10% and 400 mL of 20%

OpenStudy (mertsj):

Yep

OpenStudy (anonymous):

ok